The sum of a number and 12 times its reciprocal is -7. Find the numbers.
Answer by algebrapro18(206)

first create an equation representing the situation

x+12/x=-7

now multiply both sides of the equation by x to clear the fraction

x^2+12 = -7x

now solve

x^2+7x+12=0
(x+3)(x+4)

x+3=0 x+4=0

x= -3 and -4

checking

-3-12/3=-7
-3+-4=-7
-7=-7

-4-12/-4=-7
-4+-3=-7
-7=-7